[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Mon Jun 26, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Mon Jun 26 10:45:11 UTC 2017


TAFB Shift Log for Mon Jun 26 2017 06:45:04 GMT-0400 (EDT)

Shift: Mid-Shift

Pacific Forecaster: A Levine

Atlantic Forecaster: S Stripling

Surface Analyst: N Ramos

Technical Issues Summary:

GFE/NCP issues: No GOES-16 imagery available in AWIPS2 in any sector or  
channel. Called NCF, Tyler, to investigate, but no ticket # given.

N-AWIPS issues: No GOES-16 Imagery available since 20Z both in TAFB and  
HSU. Called SDM to inform and he was aware and doing some investigation.

Other issues: Situational Awareness PC is making loud noise behind the Atlc  
desk. Sounds like the fan. Turned it off overnight and the monitors as well.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170626/210176ba/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list